OpenStudy (anonymous):

int(secxtanx+tan^2x)

OpenStudy (xapproachesinfinity):

hmm \[\int\limits \sec x \tan x+ \tan^2 x dx\] do \[\int\limits\sec x \tan x dx +\int\limits \tan ^2x dx\]

OpenStudy (xapproachesinfinity):

the first part is something that you already know the second part you need to bring forth an identity